# SemantiCodec: An Ultra Low Bitrate Semantic Audio Codec for General Sound

### v1

@@ III-B Semantic Clustering
To perform semantic quantization on the AudioMAE feature vectors ${\bm{y}}_{i}$, we utilize k-means clustering [24], a widely used technique for partitioning a dataset into clusters in which each data point belongs to the cluster with the nearest mean. Our preliminary experiment indicates that the diverse acoustic characteristics of different audio types could lead to suboptimal outcomes when a single k-means clustering model is applied to a varied audio dataset. For instance, speech may necessitate finer granularity in clustering due to its rich semantic content, unlike more homogeneous sounds such as wind or church bells. We employ an ensemble clustering approach to overcome these issues and enhance semantic clustering accuracy, motivated by the cluster ensemble approach used in HuBERT [30]. This involves training distinct k-means models for three specific audio categories: speech, music, and general sounds. The codebooks derived from these domains are then merged, forming an ensembled codebook that accommodates the distinct acoustic features of each audio category, thereby optimising the clustering process for each unique audio type.
@@ III-C SemantiCodec Encoder
As introduced in Section III-A, the encoding process of the SemantiCodec includes taking the stacked AudioMAE feature ${\bm{Y}}$ as input and calculating the token ${\bm{c}}=[{\bm{c}}_{\text{s}},{\bm{c}}_{\text{a}}]$ and the latent feature ${\bm{E}}=[{\bm{E}}_{\text{s}},{\bm{E}}_{\text{a}}]$. This part introduces the derivation of ${\bm{c}}$ and ${\bm{E}}$.
@@
$\small c_{\text{s}}(i)=\operatorname*{arg\,min}_{j\in\{1,\ldots,N_{\text{s}}\}}\|{\bm{y}}_{i}-{\bm{e}}_{j}\|^{2},{\bm{E}}_{\text{s}}(i)={\bm{e}}_{c_{\text{s}}(i)}\quad i\in\left\{1,2,\ldots,\frac{L}{K}\right\},$ (1)
where $c_{\text{s}}(i)$ denotes the index of the closest centroid in the semantic codebook $\mathbb{E}_{\text{s}}$ to the feature vector ${\bm{y}}_{i}$, and ${\bm{E}}_{\text{s}}(i)$ is the centroid vector. This quantization step effectively maps each high-dimensional feature vector ${\bm{y}}_{i}$ into a discrete semantic token $c_{\text{s}}(i)$ and a corresponding quantized semantic feature ${\bm{E}}_{\text{s}}(i)$.
While the quantized AudioMAE feature ${\bm{E}}_{\text{s}}$ encapsulates rich semantic information, our experiments indicate that using ${\bm{E}}_{\text{s}}$ alone as the condition for audio reconstruction leads to sub-optimal quality (see Table I), such as unintelligible speech. To address this, we introduce an additional acoustic encoder module with vector quantization to capture discrete acoustic detail-oriented representations. The input feature we used for acoustic quantization is calculated by an acoustic encoder $\mathcal{F}_{\Phi}(\cdot)$, which takes the stack of the AudioMAE feature before and after semantic quantization as input and outputs the acoustic feature, given by
@@
where $\Phi$ denotes the trainable parameter in the acoustic encoder. Specifically, we employ a bi-directional long short-term memory (BiLSTM) based network [51] as the acoustic encoder.
@@
$\small c_{\text{a}}(i)=\operatorname*{arg\,min}_{j\in\{1,\ldots,N_{\text{a}}\}}\|{\bm{y}}_{a,i}-{\bm{e}}_{j}\|^{2},{\bm{E}}_{\text{a}}(i)={\bm{e}}_{c_{\text{a}}(i)}\quad i\in\{1,2,\ldots,\frac{L}{K}\},$ (3)
where $c_{\text{a}}(i)$ identifies the index of the nearest centroid within the acoustic codebook $\mathbb{E}_{\text{a}}$ for the feature vector ${\bm{y}}_{a,i}$, and ${\bm{E}}_{\text{a}}(i)$ denotes the quantized vector. To encourage the acoustic encoder to produce representations that closely match the codebook entry and stabilise training, we employ a commitment loss [5], given by
@@
Following the success of HiFi-Codec [8], our approach utilizes an exponential moving average (EMA) mechanism for codebook update. The final tokens ${\bm{c}}$ and the representation ${\bm{E}}$ are the concatenation of output from both the semantic quantization layer and acoustic quantization layer, given by
@@
Note that the dual-layer vector quantization architecture (semantic quantization and acoustic quantization) in SemantiCodec is different from residual vector quantization approach [4, 6], where the second layer aims to refine the codec encoder output directly. Instead, our approach allows the output of the second layer to be freely optimized in conjunction with the diffusion decoder for end-to-end training.
@@ III-D Latent Diffusion Model for Reconstruction
Following the success of latent diffusion models (LDMs) for conditional audio generation [47, 25, 48], we employ an LDM as the decoder to reconstruct the original audio $x$. The LDM models the data distribution in a latent space constructed from a VAE, which is directly adopted from the VAE used in AudioLDM 2 [49]. Compared with the original diffusion model [52], the high-dimensional spectrogram ${\bm{X}}$ is compressed to the low-dimensional latent $\mathbf{z}_{0}$ in the LDM to significantly alleviate the computations. Then a diffusion model is trained to gradually generate $\mathbf{z}_{0}$ from the standard Gaussian noise. The forward diffusion process comprises a series of $N$ Markov transition steps that gradually transform $\mathbf{z}_{0}$ into a Gaussian distribution by noise injection. The forward step $n-1$ is defined as:
@@
where $\alpha_{n}=1-\beta_{n},\bar{\alpha}_{n}=\prod_{n=1}^{n}\alpha_{n}$ and $\epsilon\sim\mathcal{N}(0,I)$. With enough diffusion steps $N$, $q(\mathbf{z}_{n})$ will approximate a standard Gaussian distribution $\mathcal{N}(0,I)$. The LDM is trained to model the reverse probability $p_{\theta}(\mathbf{z}_{n-1}|\mathbf{z}_{n},{\bm{E}})$ conditioned on the SemantiCodec encoder output ${\bm{E}}$.
@@
We adopt the DDIM sampler [55] for inference. Finally, the audio $\hat{{\bm{x}}}$ is reconstructed via the pretrained VAE decoder and a vocoder. The vocoder is a HiFi-GAN-based architecture [56] and is directly adopted from the pretrained AudioLDM 2. Our LDM adopts the Transformer-UNet (T-UNet) introduced in AudioLDM 2 [49], which incorporates self-attention and cross-attention Transformer blocks between convolutional blocks, significantly enhancing the model capacity to complex audio patterns. However, our experiments suggest that the impact of T-UNet parameter numbers, such as $75$ million or $346$ million, on the quality of audio reconstruction is relatively minor compared to their role in text-to-audio generation tasks. Therefore, we adopt a T-UNet with reduced parameter size compared with AudioLDM 2.
To achieve potentially better quality in reconstruction, we adopt classifier-free guidance [57, 58] (CFG), a common approach to guiding the audio generation by the diffusion models. The condition ${\bm{E}}$ in Equation 9 is randomly discarded with a certain probability during training so that both conditional generation models $v_{\theta}(\mathbf{z}_{n},n,{\bm{E}})$ and unconditional generation models $v_{\theta}(\mathbf{z}_{n},n)$ are optimized in a multi-task paradigm. During sampling, the original $v_{\theta}(\mathbf{z}_{n},n,{\bm{E}})$ is replaced by the weighted combination of velocities predicted by conditional and unconditional models:
@@ IV-A1 Training Datasets
Our model training is supported by various datasets, which can be generally classified into three categories: speech, music, and general sounds. For speech, we utilize the GigaSpeech (GGS) [59] dataset, a comprehensive English speech recognition corpus with around $10,000$ hours of transcribed audio, and the speech dataset collected by VoiceFixer [60] on OpenSLR [^1] https://openslr.org/, featuring a multi-lingual speech dataset with $186,514$ short audio clips. The music category includes the Million Song Dataset (MSD)[61], which provides a vast collection of $510,000$ music tracks with metadata. We also adopt datasets including MedleyDB [62], and MUSDB18 [63] training subset, mostly used for music source separation [64]. For general audio sounds, we engaged with AudioSet (AS)[65], the largest classification dataset offering two million ten-second audio clips across $527$ categories, WavCaps[66] that includes ChatGPT-assisted weakly-labeled audio captions for $403,050$ clips, and VGGSound (VS)[67], a substantial audio-visual dataset with around $190,000$ videos from which we only utilized the audio data. All the audio data are resampled to $16$kHz during training and evaluation. Our ablation studies in Section V-C2 and Section V-C3 utilize $10$% of the training data to speed up the model training.
@@ IV-A2 Reconstruction Performance Evaluation
To assess the reconstruction capabilities of our audio codec, we carefully select and evaluate three distinct categories of datasets similar to building the training set. Within the speech category, we carefully choose a subset of the LibriTTS clean test set [68], selecting $300$ speech utterances randomly. These utterances, each lasting between $8$ and $10$ seconds, come with detailed transcription annotations, offering a rich basis for evaluating speech reconstruction accuracy. We randomly select $500$ segments from the AudioSet evaluation set for general audio evaluation, ensuring a diverse representation of ambient sounds, effects, and non-musical content. Our music data evaluation leverages the MUSDB18 test set [63], comprising $50$ songs with isolated tracks for vocals, drums, bass, and other elements. From each of the four tracks and their mixture for every song, we randomly select a $10$-second segment, ensuring it is non-silent, to form our evaluation set. Our evaluation dataset encompasses $1050$ audio samples, achieving a relatively balanced distribution across speech, music, and general sounds. Our MUSHRA (MUltiple Stimuli with Hidden Reference and Anchor) test is performed on the same evaluation dataset while only $10\%$ of the data is used to save effort for subjective evaluation. Our evaluation set and metrics implementations are publicly available [^2] https://zenodo.org/records/11047204.
@@ IV-A3 Semantic Information Evaluation
To assess the richness of semantic information captured by audio codec representations, we employ a diverse array of datasets following a subset of the HEAR evaluation benchmark [22]. We choose the tasks oriented towards clip-level audio content classification, eliminating tasks like Gunshot Triangulation, which aims to recognize recording devices. Also, for convenience, we choose the datasets where the duration of each audio clip is shorter or close to $10$ seconds. The evaluation datasets we employed include the following: (i) NSynth Pitch (NSPitch) [69] is utilized to evaluate the ability to recognize musical pitches, a fundamental aspect of music theory and auditory perception. (ii) ESC-50 [70] encompasses a broad array of environmental sounds such as rainstorms and animal calls, testing the versatility of the codec feature towards general sound effects. (iii) LibriCount (LbCount) [71] is the dataset used to determine the number of speakers in an audio clip, combining speech detection with the differentiation of speakers within complex auditory scenes. (iv) CREMA-D (CRM-D) [72] focuses on speech emotion recognition, requiring models to classify a wide spectrum of emotional states conveyed through speech. (v) Vocal Imitations (VoImit) [73] dataset examines the ability to classify non-verbal human vocal imitations of various sounds. (vi) Speech Commands (SC) [74] dataset tests the recognition of specific spoken commands, emphasizing speech clarity and command accuracy. These audio classification datasets collectively offer a comprehensive evaluation framework, spanning musical notes, environmental sounds, speech nuances, and non-verbal vocalizations, to thoroughly assess the semantic capabilities of audio codec representations.
@@ IV-B Baselines
We employ several state-of-the-art neural audio codecs as baselines for comparison, including Encodec (EC), Descript Codec (DAC), and HiFi-Codec (HC), which have demonstrated success in the domain of general sound. The open-sourced Encodec [^3] https://github.com/facebookresearch/encodec is trained on a variety of $24$kHz sampling rate audio data and can compress audio to $1.5$, $3.0$, $6.0$, $12.0$, and $24.0$ kbps. Since SemantiCodec operates at a relatively low bitrate, we only compare Encodec with $1.5$, $3.0$, and $6.0$ kbps settings. For DAC, we adopt the open-sourced $6.0$kbps model that operates on $16$kHz as one of the baselines. Since DAC does not have checkpoint open-sourced for a lower bitrate, we train another three variants of the DAC with the same training data as SemantiCodec following the setting described in their paper [7]. The three DAC settings have a bit rate of $0.47$, $0.78$, and $1.41$ kbps, which are comparable to the three variants of SemantiCodec. For HiFi-Codec, we adopt the $2.0$ kbps checkpoint [^4] https://github.com/yangdongchao/AcademiCodec.

### v2

@@ III-B Semantic Clustering
To perform semantic quantization on the AudioMAE feature vectors ${\bm{y}}_{i}$, we utilize k-means clustering [24], a widely used technique for partitioning a dataset into clusters in which each data point belongs to the cluster with the nearest mean. Our preliminary experiment indicates that the diverse acoustic characteristics of different audio types could lead to suboptimal outcomes when a single k-means clustering model is applied to a varied audio dataset. For instance, speech may necessitate finer granularity in clustering due to its rich semantic content, unlike more homogeneous sounds such as wind or church bells. We employ an ensemble clustering approach to overcome these issues and enhance semantic clustering accuracy, motivated by the cluster ensemble approach used in HuBERT [30]. This involves training distinct k-means models for three specific audio categories: speech, music, and general sounds. The codebooks (i.e., k-means cluster centroids) derived from these domains are then combined together, forming a new ensembled codebook that accommodates the distinct acoustic features of each audio category.
@@ III-C SemantiCodec Encoder
As introduced in Section III-A, the encoding process of the SemantiCodec includes taking the stacked AudioMAE feature ${\bm{Y}}$ as input and calculating the tokens ${\bm{c}}=[{\bm{c}}_{\text{s}},{\bm{c}}_{\text{a}}]$ and the latent features ${\bm{E}}=[{\bm{E}}_{\text{s}},{\bm{E}}_{\text{a}}]$. This part introduces the derivation of ${\bm{c}}$ and ${\bm{E}}$.
III-C1 Semantic Encoder
@@
$c_{\text{s}}(i)=\operatorname*{arg\,min}_{j\in\{1,\ldots,N_{\text{s}}\}}\|{\bm{y}}_{i}-{\bm{e}}_{j}\|^{2},\quad{\bm{E}}_{\text{s}}(i)={\bm{e}}_{c_{\text{s}}(i)},$ (1)
where $i\in\{0,1,\ldots,\frac{L}{K}-1\}$, $c_{\text{s}}(i)$ denotes the index of the closest centroid in the semantic codebook $\mathbb{E}_{\text{s}}$ to the feature vector ${\bm{y}}_{i}$, and ${\bm{E}}_{\text{s}}(i)$ is the centroid vector. This quantization step effectively maps each high-dimensional feature vector ${\bm{y}}_{i}$ into a discrete semantic token $c_{\text{s}}(i)$ and corresponding quantized semantic feature ${\bm{E}}_{\text{s}}(i)$.
III-C2 Acoustic Encoder
While the quantized AudioMAE feature ${\bm{E}}_{\text{s}}$ encapsulates rich semantic information, our experiments indicate that using ${\bm{E}}_{\text{s}}$ alone as the condition for audio reconstruction leads to sub-optimal quality (see Table II), such as unintelligible speech. To address this, we introduce an additional acoustic encoder module with vector quantization to capture discrete acoustic detail-oriented representations. The input feature we used for acoustic quantization is calculated by an acoustic encoder $\mathcal{F}_{\Phi}(\cdot)$, which takes the stack of the AudioMAE feature before and after semantic quantization as input and outputs the acoustic feature, given by
@@ III-C2 Acoustic Encoder
where $\Phi$ denotes the trainable parameter in the acoustic encoder. Specifically, we employ a bi-directional long short-term memory (BiLSTM) based network [51] as the acoustic encoder. We use both ${\bm{Y}}$ and ${\bm{E}}_{\text{s}}$ as inputs because ${\bm{E}}_{\text{s}}$ undergoes information loss during semantic quantization, while ${\bm{Y}}$ retains most of the necessary information for audio reconstruction. By providing both ${\bm{Y}}$ and ${\bm{E}}_{\text{s}}$, the second acoustic quantization layer can effectively compensate for the information lost during semantic quantization.
@@
$c_{\text{a}}(i)=\operatorname*{arg\,min}_{j\in\{1,\ldots,N_{\text{a}}\}}\|{\bm{y}}_{a,i}-{\bm{e}}_{j}\|^{2},\quad{\bm{E}}_{\text{a}}(i)={\bm{e}}_{c_{\text{a}}(i)},$ (3)
where $i\in\{0,1,\ldots,\frac{L}{K}-1\}$, $c_{\text{a}}(i)$ identifies the index of the nearest centroid within the acoustic codebook $\mathbb{E}_{\text{a}}$ for the feature vector ${\bm{y}}_{a,i}$, and ${\bm{E}}_{\text{a}}(i)$ denotes the quantized vector. To encourage the acoustic encoder to produce representations that closely match the codebook entries and stabilise training, we employ a commitment loss [5], given by
@@
Following the success of HiFi-Codec [8], our approach utilizes an exponential moving average (EMA) mechanism for codebook update. The final tokens ${\bm{c}}$ and the representation ${\bm{E}}$ are the concatenation of outputs from both the semantic quantization layer and acoustic quantization layer, given by
@@
Note that the dual-layer vector quantization architecture (semantic quantization and acoustic quantization) in SemantiCodec is different from the residual vector quantization (RVQ) approach [4, 6], where the second layer aims to refine the residual from the first VQ layer. Although both SemantiCodec and RVQ rely on multiple layers of vector quantization, SemantiCodec does not explicitly calculate residuals between each layer. Instead, SemantiCodec concatenates the outputs of each VQ layer to form the output of the encoder.
Our experiments in Section V demonstrate that the dual-layer vector quantization in SemantiCodec effectively decouple the semantic and acoustic information. Specifically, removing the acoustic quantization layer significantly degrades the reconstruction quality (Table II), while having only a minor impact on the semantic information in the latent space (Table III).
@@ III-D Latent Diffusion Model for Reconstruction
Following the success of latent diffusion models (LDMs) for conditional audio generation [47, 25, 48], we employ an LDM as the decoder to reconstruct the original audio $x$. The LDM models the data distribution in a latent space constructed from a VAE, which is directly adopted from the VAE used in AudioLDM 2 [49]. Note that the VAE encoder is only used during SemantiCodec training, and is discarded during SemantiCodec inference. Compared with the original diffusion model [52], the high-dimensional spectrogram ${\bm{X}}$ is compressed to the low-dimensional latent $\mathbf{z}_{0}$ in the LDM to significantly alleviate the computations. Then a diffusion model is trained to gradually generate $\mathbf{z}_{0}$ from Gaussian noise. The forward diffusion process comprises a series of $N$ Markov transition steps that gradually transform $\mathbf{z}_{0}$ into a Gaussian distribution by noise injection. The forward step $n-1$ is defined as:
@@
where $\alpha_{n}=1-\beta_{n},\bar{\alpha}_{n}=\prod_{n=1}^{n}\alpha_{n}$ and $\epsilon\sim\mathcal{N}(0,I)$. With enough diffusion steps $N$, $q(\mathbf{z}_{n})$ will approximate a standard Gaussian distribution $\mathcal{N}(0,I)$. The LDM is trained to model the reverse probability $p_{\theta}(\mathbf{z}_{n-1}|\mathbf{z}_{n},{\bm{E}})$ conditioned on the SemantiCodec encoder output ${\bm{E}}$. Although ${\bm{E}}_{\text{a}}$ may already contain part of the information from ${\bm{E}}_{\text{s}}$, we feed both into the diffusion model as conditions to enable both quantization layers to collaboratively capture as much information from ${\bm{Y}}$ as possible.
@@
For model inference, we adopt the Denoising Diffusion Implicit Models (DDIM) sampler [55], which uses non-Markovian diffusion processes for sampling, allowing for computationally efficient sampling while maintaining high generation quality. Finally, the audio $\hat{{\bm{x}}}$ is reconstructed via the pretrained VAE decoder and a vocoder. The vocoder is a HiFi-GAN-based architecture [56] and is directly adopted from the pretrained AudioLDM 2. Our LDM adopts the Transformer-UNet (T-UNet) introduced in AudioLDM 2 [49], which incorporates self-attention and cross-attention Transformer blocks between convolutional blocks, significantly enhancing the model capacity for complex audio patterns. However, our experiments suggest that the impact of T-UNet parameter numbers, such as $75$ million or $346$ million, on the quality of audio reconstruction is relatively minor compared to their role in text-to-audio generation tasks. Therefore, we adopt a T-UNet with reduced parameter size compared with AudioLDM 2.
To achieve potentially better quality in reconstruction, we adopt classifier-free guidance (CFG) [57, 58], a common approach to guiding the audio generation in diffusion models. The condition ${\bm{E}}$ in Equation (9) is randomly discarded with a certain probability during training so that both conditional generation models $v_{\theta}(\mathbf{z}_{n},n,{\bm{E}})$ and unconditional generation models $v_{\theta}(\mathbf{z}_{n},n)$ are optimized in a multi-task paradigm. During sampling, the original $v_{\theta}(\mathbf{z}_{n},n,{\bm{E}})$ is complemented by the weighted combination of velocities predicted by conditional and unconditional models:
@@ IV-A1 Training Datasets
Our model training is supported by various datasets, which can be generally classified into three categories: speech, music, and general sounds. For speech, we utilize the GigaSpeech (GGS) [59] dataset, a comprehensive English speech recognition corpus with around $10,000$ hours of transcribed audio, and the speech dataset collected by VoiceFixer [60] on OpenSLR [^1] https://openslr.org/, featuring a multi-lingual speech dataset with $186,514$ short audio clips. The music category includes the Million Song Dataset (MSD)[61], which provides a vast collection of $510,000$ music tracks with metadata. We also adopt datasets including MedleyDB [62], and MUSDB18 [63] training subset, mostly used for music source separation [64]. For general sounds, we engaged with AudioSet (AS)[65], the largest classification dataset offering two million ten-second audio clips across $527$ categories, WavCaps[66] that includes ChatGPT-assisted weakly-labeled audio captions for $403,050$ clips, and VGGSound (VS)[67], a substantial audio-visual dataset with around $190,000$ videos from which we only utilized the audio data. All the audio data are in $16$ bits and resampled to $16$ kHz during training and evaluation. Therefore, all SemantiCodec configurations presented in this work operate at a $16$ kHz sampling rate. We plan to explore higher sampling rate versions of SemantiCodec in future research. Our ablation studies in Section V-C2 and Section V-C3 utilize only $10$% of the training data to speed up the model training.
@@ IV-A2 Reconstruction Performance Evaluation
To assess the reconstruction capabilities of our audio codec, we carefully select and evaluate three distinct categories of datasets similar to building the training set. Within the speech category, we choose a subset of the LibriTTS clean test set [68], selecting $300$ speech utterances randomly. These utterances, each lasting between $8$ and $10$ seconds, come with detailed transcription annotations, offering a rich basis for evaluating speech reconstruction accuracy. We randomly select $500$ segments from the AudioSet evaluation set for general sound evaluation, ensuring a diverse representation of ambient sounds, effects, and non-musical content. Our music data evaluation leverages the MUSDB18 test set [63], comprising $50$ songs with isolated tracks for vocals, drums, bass, and other elements. From each of the four tracks and their mixture for every song, we randomly select a $10$-second segment, ensuring it is non-silent, to form our evaluation set. Our evaluation dataset encompasses $1050$ audio samples, achieving a relatively balanced distribution across speech, music, and general sounds. Our MUSHRA (MUltiple Stimuli with Hidden Reference and Anchor) test [69] is performed on the same evaluation dataset while only $10\%$ of the data is used to save effort for subjective evaluation. Our evaluation set and metrics implementations are publicly available [^2] https://zenodo.org/records/11047204.
@@ IV-A3 Semantic Information Evaluation
To assess the richness of semantic information captured by audio codec representations, we employ a diverse array of datasets following a subset of the Holistic Evaluation of Audio Representations (HEAR) benchmark [22]. We choose the tasks oriented towards audio classification [70], eliminating tasks like Gunshot Triangulation, which aims to recognize recording devices. Also, for convenience, we choose the datasets where the duration of each audio clip is shorter or close to $10$ seconds. The evaluation datasets we employed include the following: (i) NSynth Pitch (NSPitch) [71] is utilized to evaluate the ability to recognize musical pitches, a fundamental aspect of music theory and auditory perception. (ii) ESC-50 [72] encompasses a broad array of environmental sounds such as rainstorms and animal calls, testing the versatility of the codec features towards general sound effects. (iii) LibriCount (LbCount) [73] is the dataset used to determine the number of speakers in an audio clip, combining speech detection with the differentiation of speakers within complex auditory scenes. (iv) CREMA-D (CRM-D) [74] focuses on speech emotion recognition, requiring models to classify a wide spectrum of emotional states conveyed through speech. (v) Vocal Imitations (VoImit) [75] dataset examines the ability to classify non-verbal human vocal imitations of various sounds. (vi) Speech Commands (SC) [76] dataset tests the recognition of specific spoken commands, emphasizing speech clarity and command accuracy. These audio classification datasets collectively offer a comprehensive evaluation framework, spanning musical notes, environmental sounds, speech nuances, and non-verbal vocalizations, to thoroughly assess the semantic capabilities of audio codec representations.
@@ IV-B Baselines
We employ several state-of-the-art neural audio codecs as baselines for comparison, including Encodec (EC), Descript Codec (DAC), and HiFi-Codec (HC), which have demonstrated success in the domain of general audio. The open-sourced Encodec [^3] https://github.com/facebookresearch/encodec is trained on a variety of audio datasets sampled at $24$ kHz and can compress audio to $1.5$, $3.0$, $6.0$, $12.0$, and $24.0$ kbps. Since SemantiCodec operates at a relatively low bitrate, we only include Encodec at $1.5$, $3.0$ and $6.0$ kbps for comparisons. For DAC, we adopt the open-sourced $6.0$ kbps model that operates on $16$ kHz as one of the baselines. As DAC does not have checkpoints open-sourced for lower bitrates, we train three new versions of DAC on the same training data as SemantiCodec following the setting described in their paper [7]. The three DAC settings have a bit rate of $0.47$, $0.78$, and $1.41$ kbps, which are comparable to the three variants of SemantiCodec. For HiFi-Codec, we adopt the $2.0$ kbps checkpoint [^4] https://github.com/yangdongchao/AcademiCodec. The baseline models we used, including Encodec, Descript Codec, and HiFi-Codec, contain approximately $23$, $74$, and $63$ million parameters, respectively.